Circular No. 7135

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999bg IN IC 758 W. Li, University of California at Berkeley, reports the discovery of an

apparent supernova by the Lick Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1999bg was found and confirmed on unfiltered CCD frames taken on Mar. 28.3 and 29.3 UT (mag ∼ 15.5 on both dates). The new object is located at α = 12^{h}04^{m}07^{s}.30, δ = +62^{o}30′01′′.2 (equinox 2000.0), which is ∼ 33′′ west and 20′′ south of the nucleus of IC 758. A KAIT image of the same field on Feb. 28.4 (limiting mag ∼ 18.0) showed nothing at the position of SN 1999bg.

SUPERNOVA 1999bh IN NGC 3435 Li also reports the discovery by LOSS of another apparent supernova

in an unfiltered image taken on Mar. 29.2 UT (mag ∼ 16.8). SN 1999bh was confirmed at mag 17.5 on an earlier image taken on Mar. 24.2 and is located at α = 10^{h}54^{m}46^{s}.97, δ = +61^{o}17′20′′.0 (equinox 2000.0), which is ∼ 10′′ west and 3′′ south of the nucleus of NGC 3435. A KAIT image of the same field on Mar. 17.2 (limiting mag ∼ 19.0) already showed a hint of SN 1999bh, and a KAIT image on Mar. 12.2 (limiting mag ∼ 19.0) showed nothing at the position of the supernova.

COMET C/1995 O1 (HALE-BOPP)

Ephemeris extension to IAUC 7083 (elements MPC 32410):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 2 4^{h}53.^{m}44 −65^{o}55.^{′}6 8.119 8.067 83.^{o}5 7.^{o}1 11.6
12 4 58.99 −65 06.0 8.193 8.147 83.8 7.0 11.7
22 5 05.65 −64 23.4 8.266 8.226 84.2 7.0 11.7
May 2 5 13.20 −63 48.7 8.338 8.305 84.6 6.9 11.8
12 5 21.43 −63 22.1 8.409 8.383 85.1 6.9 11.9
22 5 30.20 −63 04.2 8.481 8.461 85.5 6.8 11.9
June 1 5 39.33 −62 55.2 8.552 8.539 85.9 6.8 12.0
11 5 48.70 −62 55.0 8.624 8.616 86.2 6.8 12.0
21 5 58.19 −63 03.7 8.696 8.694 86.5 6.7 12.1
July 1 6 07.67 −63 21.1 8.770 8.770 86.7 6.6 12.1
11 6 17.03 −63 46.9 8.845 8.847 86.8 6.6 12.2
21 6 26.13 −64 20.6 8.921 8.923 86.9 6.5 12.3
1999 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green